Two resistors have resistances R(smaller) and R(larger), where R(smaller) < R(larger). When the resistors are connected in series to a 12.0-V battery, the current from the battery is 1.77 A. When the resistors are connected in parallel to the battery, the total current from the battery is 11.3 A. Determine the two resistances.
Answer #1

When connected in series

R1 + R2 = V/I = 12/1.77

R1 + R2 = 6.78

R1 = 6.78 - R2 .. (I)

When connected in parallel

R1 * R2/ ( R1 + R2) = 12/ 11.3

From (I)

(6.78 - R2 ) * R2 = 1.06 * ( 6.78)

6.78R2 - R2^2 = 7.19

R2^2 - 6.78R2 + 7.19 = 0

R2 = 5.46 ohm (Ans)

From (I)

R1 = 6.78 - 5.46 = 1.32 ohm (Ans)

Comment in case any doubt please rate my answer ....